OpenStudy (anonymous):

(sinx+cosx)^2-1=sin2x

OpenStudy (solomonzelman):

Are you aware of the following rule? \(\color{#000000 }{ \displaystyle (a+b)^2=a^2+2ab+b^2 }\)

OpenStudy (solomonzelman):

Expand the left side, please (using this property).

OpenStudy (anonymous):

THANK YOU

OpenStudy (solomonzelman):

We aren't done are we?

OpenStudy (anonymous):

We're doing the work now. We needed the help getting started. What we've done now is use the Pythagorean Identity to cancel out the sin^2x and cos^2x given when we foiled and can cancel out the minus one. The sin2x can be changed to a double and identity of 2sinxcosx which is what is left once everything else on the left is canceled. Thank you!

OpenStudy (solomonzelman):

Wow, that was fabulous!! We rarely have users that are actually capable of doing their work; good job!

OpenStudy (solomonzelman):

:)